OpenStudy (anonymous):

A $4,000 loan, taken now, with a simple interest rate of 7% per year, will require a total repayment of $5,960. At what time t will the loan mature?

OpenStudy (anonymous):

A = p + prt is the formula for the total you will receive after t years, r = rate of simple interest, and p = principal (amount you put into the bank) so A = 5960 5960 = 4000 + 4000(.07) (t) Let's solve for t, the # of years the money was in the bank 5960 = 4000 + 280t 1960 = 280t t = 7 The money was in the bank for 7 years.